# This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible

### v3

@@ IV Experiments
This section presents the empirical results of our extensive experiments. We selected five representative models and three scaling techniques, along with a baseline without scaling, as this subset already allows us to draw meaningful conclusions and observe distinctions among models. The complete results are available in the Appendix A.
IV-A Impact on Validation Metrics
TABLE III: Accuracy results by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Breast Cancer Wisconsin Diagnostic KNN 0.9591 0.9766 0.9591 0.9649
LGBM 0.9474 0.9474 0.9591 0.9591
MLP 0.9649 0.9766 0.9766 0.9708
RF 0.9708 0.9708 0.9708 0.9708
SVM 0.9240 0.9825 0.9766 0.9825
Dry Bean KNN 0.7113 0.9141 0.9216 0.9190
LGBM 0.9275 0.9275 0.9263 0.9275
MLP 0.2980 0.9101 0.9327 0.9314
RF 0.9238 0.9226 0.9226 0.9226
SVM 0.5803 0.9109 0.9263 0.9268
Glass Identification KNN 0.5846 0.6615 0.6308 0.6308
LGBM 0.8154 0.8154 0.8000 0.8308
MLP 0.7077 0.7231 0.6923 0.6769
RF 0.7538 0.7538 0.7692 0.7846
SVM 0.6769 0.5077 0.6615 0.6462
Heart Disease KNN 0.4945 0.5495 0.5714 0.5275
LGBM 0.5275 0.5275 0.5275 0.5385
MLP 0.3516 0.5495 0.5055 0.5385
RF 0.5604 0.5604 0.5604 0.5604
SVM 0.5055 0.5934 0.6154 0.5934
Iris KNN 1.0000 1.0000 1.0000 0.9556
LGBM 1.0000 1.0000 1.0000 1.0000
MLP 1.0000 1.0000 1.0000 1.0000
RF 1.0000 1.0000 1.0000 1.0000
SVM 1.0000 1.0000 0.9778 0.9778
Letter Recognition KNN 0.9493 0.9480 0.9405 0.9158
LGBM 0.9640 0.9640 0.9637 0.9637
MLP 0.9367 0.9280 0.9502 0.9545
RF 0.9577 0.9577 0.9570 0.9580
SVM 0.8135 0.8208 0.8488 0.8488
Magic Gamma Telescope KNN 0.8098 0.8254 0.8340 0.8340
LGBM 0.8803 0.8803 0.8792 0.8810
MLP 0.8170 0.8677 0.8717 0.8777
RF 0.8808 0.8808 0.8808 0.8808
SVM 0.2976 0.5158 0.4341 0.3212
Rice Cammeo And Osmancik KNN 0.8775 0.9204 0.9143 0.9064
LGBM 0.9213 0.9213 0.9178 0.9160
MLP 0.5468 0.9335 0.9318 0.9309
RF 0.9265 0.9265 0.9265 0.9265
SVM 0.9248 0.9309 0.9309 0.9274
Wine KNN 0.7407 0.9444 0.9630 0.9444
LGBM 0.9815 0.9815 0.9815 0.9815
MLP 0.9815 1.0000 0.9815 0.9815
RF 1.0000 1.0000 1.0000 1.0000
SVM 0.5926 1.0000 0.9815 0.9815
TABLE IV: $R^{2}$ score by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVR under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Abalone KNN 0.5164 0.4955 0.4662 0.4552
LGBM 0.5260 0.5260 0.5256 0.5190
MLP 0.5245 0.5265 0.5578 0.5632
RF 0.5244 0.5249 0.5234 0.5241
SVR 0.5293 0.5257 0.5421 0.5398
Air Quality KNN 0.9995 0.9993 0.9994 0.9986
LGBM 0.9999 0.9999 0.9999 0.9999
MLP 0.9985 1.0000 1.0000 0.9999
RF 1.0000 1.0000 1.0000 1.0000
SVR 0.9966 0.9619 0.9269 0.9188
Appliances Energy Prediction KNN 0.1681 0.2049 0.3279 0.2929
LGBM 0.4318 0.4318 0.4334 0.4192
MLP 0.1598 0.1581 0.3144 0.2970
RF 0.5122 0.5120 0.5122 0.5125
SVR -0.1056 -0.0275 0.0154 -0.0096
Concrete Compressive Strength KNN 0.6770 0.6631 0.6714 0.7446
LGBM 0.9229 0.9229 0.9217 0.9226
MLP 0.8030 0.7468 0.8725 0.8662
RF 0.8896 0.8895 0.8891 0.8894
SVR 0.2259 0.5394 0.6093 0.6987
Forest Fires KNN -0.0115 -0.0470 -0.0447 -0.0345
LGBM -0.0246 -0.0246 -0.0188 -0.0135
MLP -0.0067 0.0057 0.0013 0.0121
RF -0.1060 -0.1101 -0.1093 -0.1058
SVR -0.0257 -0.0246 -0.0244 -0.0245
Real Estate Valuation KNN 0.6232 0.6232 0.6153 0.6348
LGBM 0.7001 0.7001 0.7120 0.7075
MLP 0.6199 0.5608 0.6436 0.6894
RF 0.7444 0.7449 0.7444 0.7444
SVR 0.4897 0.5327 0.5788 0.5872
Wine Quality KNN 0.1221 0.3283 0.3475 0.3234
LGBM 0.4557 0.4557 0.4578 0.4602
MLP 0.2500 0.3243 0.3872 0.3879
RF 0.4985 0.4991 0.4982 0.4992
SVR 0.1573 0.3185 0.3842 0.3799
TABLE V: MSE by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVR under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Abalone KNN 4.9107 5.1229 5.4200 5.5324
LGBM 4.8137 4.8137 4.8170 4.8846
MLP 4.8285 4.8087 4.4900 4.4356
RF 4.8295 4.8244 4.8392 4.8328
SVR 4.7801 4.8158 4.6502 4.6734
Air Quality KNN 0.9109 1.2492 1.0231 2.3434
LGBM 0.1129 0.1129 0.1298 0.1291
MLP 2.6090 0.0278 0.0110 0.1181
RF 0.0131 0.0131 0.0131 0.0131
SVR 5.7798 65.6306 125.8177 139.7765
Appliances Energy Prediction KNN 8570.5671 8192.0770 6924.2331 7285.3653
LGBM 5854.3508 5854.3508 5837.8269 5983.4385
MLP 8656.1894 8673.7742 7063.7610 7243.1268
RF 5025.6175 5028.0530 5025.4432 5022.6680
SVR 11390.9162 10585.9003 10144.3973 10401.7762
Concrete Compressive Strength KNN 87.4024 91.1487 88.9093 69.1034
LGBM 20.8520 20.8520 21.1791 20.9335
MLP 53.3124 68.5215 34.5083 36.1983
RF 29.8643 29.8945 30.0114 29.9231
SVR 209.4406 124.6258 105.7187 81.5215
Forest Fires KNN 8049.5193 8331.5751 8314.0348 8232.2422
LGBM 8153.9240 8153.9240 8107.4363 8065.4234
MLP 8010.9478 7912.5614 7947.6005 7861.6344
RF 8801.7409 8833.7029 8827.5833 8799.6126
SVR 8162.5768 8154.0176 8152.2006 8152.5913
Real Estate Valuation KNN 63.0027 63.0182 64.3252 61.0770
LGBM 50.1547 50.1547 48.1590 48.9211
MLP 63.5585 73.4388 59.6005 51.9328
RF 42.7505 42.6657 42.7479 42.7405
SVR 85.3424 78.1453 70.4354 69.0369
Wine Quality KNN 0.6406 0.4901 0.4761 0.4937
LGBM 0.3971 0.3971 0.3956 0.3939
MLP 0.5472 0.4930 0.4471 0.4466
RF 0.3659 0.3655 0.3661 0.3654
SVR 0.6149 0.4973 0.4493 0.4525
TABLE VI: MAE by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVR under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Abalone KNN 1.5673 1.6102 1.6555 1.6654
LGBM 1.5476 1.5476 1.5500 1.5632
MLP 1.6159 1.5748 1.5318 1.4958
RF 1.5590 1.5584 1.5617 1.5619
SVR 1.5048 1.5111 1.4964 1.4990
Air Quality KNN 0.5506 0.6755 0.5862 0.8379
LGBM 0.0677 0.0677 0.0728 0.0731
MLP 1.2505 0.0909 0.0698 0.1910
RF 0.0167 0.0167 0.0166 0.0167
SVR 0.8970 1.4802 1.9293 5.4424
Appliances Energy Prediction KNN 47.7696 45.5977 39.8189 41.5828
LGBM 39.0633 39.0633 39.0969 39.3198
MLP 53.8080 54.1215 47.7318 47.7236
RF 34.2701 34.2923 34.2876 34.2838
SVR 48.9182 45.5442 43.4164 45.1344
Concrete Compressive Strength KNN 7.2301 7.0405 7.3319 6.4237
LGBM 3.0480 3.0480 3.0473 3.0197
MLP 5.9685 6.4045 4.4758 4.5392
RF 3.7512 3.7503 3.7608 3.7550
SVR 11.6674 8.9978 8.1314 7.0217
Forest Fires KNN 21.5065 20.7837 21.0229 19.7818
LGBM 24.2821 24.2821 24.1922 23.9371
MLP 21.2084 20.6192 24.5951 23.9833
RF 24.2812 24.3738 24.3496 24.3443
SVR 14.9474 14.9755 14.9655 14.9590
Real Estate Valuation KNN 5.4626 5.4435 5.6979 5.5819
LGBM 4.8106 4.8106 4.7284 4.8254
MLP 5.3601 6.1126 5.4400 5.0138
RF 4.3971 4.3951 4.4002 4.3916
SVR 6.8824 6.2845 5.9805 5.8563
Wine Quality KNN 0.6243 0.5231 0.5259 0.5349
LGBM 0.4832 0.4832 0.4829 0.4833
MLP 0.5692 0.5499 0.5187 0.5196
RF 0.4366 0.4368 0.4370 0.4362
SVR 0.6076 0.5453 0.5111 0.5142
As anticipated, one of the key findings from our classification experiments is the differential impact of feature scaling on model performance. Table III shows the accuracy results. Ensemble methods, including Random Forest and the gradient boosting family (LightGBM, CatBoost, XGBoost), demonstrated strong robustness by consistently achieving high validation performance irrespective of the preprocessing strategy or dataset. This inherent robustness offers a significant practical advantage, particularly in resource-constrained environments, since omitting the scaling step eliminates the associated memory and computational overhead. The Naive Bayes model showed similar scaling resistance, although its overall accuracy was not competitive with these top-tier ensembles. In stark contrast, the performance of Logistic Regression (LR), Support Vector Machines (SVM), K-Nearest Neighbor (KNN), TabNet, and Multi-Layer Perceptrons (MLP) was highly dependent on the choice of scaler, revealing their pronounced sensitivity to data preprocessing through significant fluctuations in performance.
A similar pattern of scaling sensitivity was observed in regression tasks when employing the regression counterparts of these classification models, as shows Tables IV, V and VI. This suggests that the underlying mathematical principles governing these model families lead to consistent behavior regarding data scaling, whether applied to classification or regression problems.
@@ IV-A Impact on Validation Metrics
IV-B Impact on Training and Inference Times
TABLE VII: Time to Train (s) in classification models by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Breast Cancer Wisconsin Diagnostic KNN 0.0002 0.0002 0.0002 0.0002
LGBM 0.0213 0.0219 0.0275 0.0256
MLP 0.1708 0.2929 0.1683 0.1824
RF 0.0706 0.0707 0.0709 0.0706
SVM 0.0016 0.0007 0.0009 0.0009
Dry Bean KNN 0.0008 0.0008 0.0008 0.0008
LGBM 0.2732 0.2759 0.2988 0.3025
MLP 0.3396 3.2012 2.4228 3.8822
RF 1.8703 1.8860 1.8872 1.8734
SVM 0.1096 0.1932 0.1560 0.1470
Glass Identification KNN 0.0002 0.0002 0.0002 0.0004
LGBM 0.0422 0.0427 0.0345 0.0410
MLP 0.2307 0.2416 0.2376 0.2372
RF 0.0453 0.0456 0.0453 0.0453
SVM 0.0010 0.0007 0.0008 0.0009
Heart Disease KNN 0.0003 0.0003 0.0003 0.0003
LGBM 0.0466 0.0450 0.0412 0.0439
MLP 0.0198 0.1105 0.3423 0.3942
RF 0.0449 0.0452 0.0451 0.0451
SVM 0.0029 0.0009 0.0017 0.0012
Iris KNN 0.0004 0.0002 0.0002 0.0003
LGBM 0.0116 0.0119 0.0130 0.0118
MLP 0.1074 0.1444 0.0947 0.1164
RF 0.0382 0.0387 0.0386 0.0385
SVM 0.0003 0.0004 0.0003 0.0003
Letter Recognition KNN 0.0010 0.0009 0.0009 0.0010
LGBM 1.1319 1.0970 1.1101 1.1380
MLP 11.1298 22.0963 11.5231 13.5357
RF 0.8526 0.8582 0.8632 0.8535
SVM 0.8484 0.7841 0.7588 0.7936
Magic Gamma Telescope KNN 0.0065 0.0065 0.0065 0.0071
LGBM 0.0469 0.0460 0.0461 0.0468
MLP 0.5592 4.8919 4.0988 3.4037
RF 2.6338 2.6361 2.6471 2.6319
SVM 0.0697 0.2742 0.2290 0.2240
Rice Cammeo And Osmancik KNN 0.0009 0.0010 0.0010 0.0010
LGBM 0.0294 0.0291 0.0309 0.0315
MLP 0.0464 0.6241 0.1954 0.2864
RF 0.2023 0.2032 0.2026 0.2034
SVM 0.0093 0.0223 0.0201 0.0199
Wine KNN 0.0003 0.0003 0.0002 0.0003
LGBM 0.0164 0.0151 0.0167 0.0154
MLP 0.2105 0.1658 0.0603 0.0696
RF 0.0406 0.0408 0.0411 0.0406
SVM 0.0008 0.0004 0.0005 0.0005
TABLE VIII: Time to Inference (s) in classification models by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Breast Cancer Wisconsin Diagnostic KNN 0.0025 0.0025 0.0025 0.0030
LGBM 0.0004 0.0004 0.0007 0.0007
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0013 0.0013 0.0013 0.0013
SVM 0.0002 0.0002 0.0001 0.0001
Dry Bean KNN 0.0529 0.0536 0.0536 0.0594
LGBM 0.0101 0.0099 0.0099 0.0100
MLP 0.0011 0.0012 0.0012 0.0012
RF 0.0180 0.0181 0.0184 0.0179
SVM 0.0255 0.1723 0.0892 0.0907
Glass Identification KNN 0.0012 0.0012 0.0012 0.0014
LGBM 0.0005 0.0006 0.0006 0.0006
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0013 0.0013 0.0013 0.0013
SVM 0.0002 0.0002 0.0002 0.0002
Heart Disease KNN 0.0016 0.0016 0.0016 0.0020
LGBM 0.0007 0.0006 0.0006 0.0006
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0014 0.0014 0.0014 0.0015
SVM 0.0003 0.0002 0.0002 0.0002
Iris KNN 0.0012 0.0010 0.0009 0.0012
LGBM 0.0004 0.0005 0.0004 0.0004
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0011 0.0011 0.0011 0.0011
SVM 0.0001 0.0001 0.0001 0.0001
Letter Recognition KNN 0.8666 0.0819 0.0821 0.0870
LGBM 0.0611 0.0606 0.0610 0.0601
MLP 0.0022 0.0022 0.0021 0.0021
RF 0.0471 0.0473 0.0472 0.0479
SVM 0.8372 1.3598 0.8540 0.8722
Magic Gamma Telescope KNN 0.1088 0.1550 0.1659 0.1635
LGBM 0.0024 0.0023 0.0024 0.0023
MLP 0.0012 0.0012 0.0012 0.0012
RF 0.0347 0.0349 0.0349 0.0341
SVM 0.0214 0.1027 0.0838 0.0856
Rice Cammeo And Osmancik KNN 0.0139 0.0148 0.0149 0.0152
LGBM 0.0007 0.0007 0.0007 0.0008
MLP 0.0003 0.0003 0.0002 0.0003
RF 0.0045 0.0044 0.0045 0.0044
SVM 0.0012 0.0083 0.0055 0.0056
Wine KNN 0.0011 0.0011 0.0011 0.0011
LGBM 0.0004 0.0004 0.0004 0.0004
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0011 0.0011 0.0011 0.0011
SVM 0.0001 0.0001 0.0001 0.0001
TABLE IX: Time to Train (s) in regression models by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Abalone KNN 0.0010 0.0010 0.0010 0.0010
LGBM 0.0286 0.0288 0.0301 0.0307
MLP 0.8414 1.1404 0.8932 1.2343
RF 0.5945 0.6035 0.6003 0.5977
SVR 0.1382 0.1383 0.1392 0.1440
Air Quality KNN 0.0032 0.0032 0.0031 0.0031
LGBM 0.0391 0.0395 0.0392 0.0396
MLP 0.5701 3.3321 2.1864 1.8880
RF 1.8482 1.8664 1.8547 1.8659
SVR 0.6315 0.5742 0.4535 0.7643
Appliances Energy Prediction KNN 0.0004 0.0004 0.0004 0.0004
LGBM 0.0522 0.0532 0.0552 0.0545
MLP 3.3408 12.6484 17.8381 16.1881
RF 18.2580 18.3987 18.3023 18.2253
SVR 3.6868 3.6063 3.6767 3.7207
Concrete Compressive Strength KNN 0.0004 0.0004 0.0004 0.0004
LGBM 0.0221 0.0232 0.0231 0.0237
MLP 0.1335 0.7281 0.8502 0.8323
RF 0.1397 0.1396 0.1390 0.1438
SVR 0.0086 0.0093 0.0087 0.0091
Forest Fires KNN 0.0003 0.0003 0.0003 0.0003
LGBM 0.0115 0.0121 0.0119 0.0118
MLP 0.2457 0.3816 0.4316 0.4330
RF 0.0997 0.1006 0.1002 0.1063
SVR 0.0025 0.0028 0.0027 0.0025
Real Estate Valuation KNN 0.0002 0.0002 0.0002 0.0002
LGBM 0.0090 0.0094 0.0094 0.0098
MLP 0.1022 0.3386 0.3599 0.3778
RF 0.0664 0.0666 0.0663 0.0689
SVR 0.0019 0.0018 0.0023 0.0017
Wine Quality KNN 0.0022 0.0022 0.0023 0.0022
LGBM 0.0314 0.0313 0.0322 0.0334
MLP 0.4713 1.3455 2.3311 1.6345
RF 1.2844 1.2891 1.2968 1.3414
SVR 0.3166 0.3181 0.3239 0.3213
TABLE X: Time to Inference (s) in regression models by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Abalone KNN 0.0034 0.0048 0.0042 0.0041
LGBM 0.0006 0.0007 0.0007 0.0007
MLP 0.0003 0.0003 0.0003 0.0003
RF 0.0110 0.0111 0.0110 0.0109
SVR 0.0652 0.0652 0.0656 0.0666
Air Quality KNN 0.0171 0.0317 0.0338 0.0381
LGBM 0.0010 0.0009 0.0009 0.0009
MLP 0.0005 0.0006 0.0006 0.0006
RF 0.0169 0.0171 0.0170 0.0171
SVR 0.2986 0.2676 0.1981 0.3599
Appliances Energy Prediction KNN 0.0222 0.0215 0.0225 0.0216
LGBM 0.0017 0.0017 0.0017 0.0017
MLP 0.0012 0.0012 0.0012 0.0012
RF 0.0694 0.0697 0.0700 0.0692
SVR 1.8656 1.8341 1.7995 1.8159
Concrete Compressive Strength KNN 0.0009 0.0010 0.0010 0.0010
LGBM 0.0004 0.0004 0.0004 0.0004
MLP 0.0001 0.0002 0.0001 0.0001
RF 0.0033 0.0034 0.0032 0.0034
SVR 0.0044 0.0043 0.0041 0.0041
Forest Fires KNN 0.0003 0.0006 0.0006 0.0005
LGBM 0.0003 0.0004 0.0003 0.0003
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0018 0.0018 0.0018 0.0018
SVR 0.0011 0.0012 0.0011 0.0013
Real Estate Valuation KNN 0.0003 0.0003 0.0003 0.0003
LGBM 0.0003 0.0003 0.0003 0.0003
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0017 0.0017 0.0017 0.0017
SVR 0.0007 0.0009 0.0007 0.0007
Wine Quality KNN 0.0048 0.0257 0.0369 0.0329
LGBM 0.0008 0.0008 0.0008 0.0009
MLP 0.0004 0.0004 0.0004 0.0005
RF 0.0143 0.0147 0.0145 0.0151
SVR 0.1501 0.1527 0.1419 0.1422
The application of different scaling techniques had a variable impact on inference times across the evaluated models. Notably, Classification and Regression Trees (CART) exhibited exceptionally robust behavior, remaining unaffected across all scaling methods and datasets. While certain Machine Learning algorithms — such as K-Nearest Neighbors (KNN), Random Forest, Support Vector Machine (SVM), and Support Vector Regressor (SVR) — showed more evident sensitivity to the choice of scaling technique, this was not the norm. For the majority of models, the preprocessing step introduced only a small, non-uniform computational overhead — as illustrated in Tables VIII and X — which may become more significant in the context of large-scale datasets or time-sensitive applications requiring real-time inference.
For training time, the effects of scaling largely mirrored those seen in the validation accuracy results, as shown in Tables VII and IX. Certain models, notably tree-based ensembles, did not derive a significant speed benefit from feature scaling, while others were more sensitive.
IV-C Results of Memory Usage (kB)
The analysis revealed a clear distinction in memory consumption between scaling techniques. As expected, applying no scaling resulted in zero additional memory consumption, only the consumption to load the data. Among the actual scaling methods, the RobustScaler, StandardScaler, Tanh Transformer, and Hyperbolic Tangent were found to be the most memory-intensive. In contrast, the MaxAbsScaler, MinMaxScaler, and Decimal Scaler consistently registered the lowest memory usage as shown in Table XI.
TABLE XI: Memory Usage (kB) per Dataset and Scaling Method
Dataset NO MA ZSN RS
Breast Cancer
Wisconsin
Diagnostic
0.1875 175.7594 176.1266 384.9979
Dry Bean 1704.2750 2448.3812 2599.3156 2388.1666
Glass
Identification
0.1875 23.1984 26.1799 51.8070
Heart Disease 33.6609 67.2016 71.7602 122.0010
Iris 0.1875 8.7828 10.5523 20.2666
Letter
Recognition
0.1875 3566.3000 3787.1906 2568.4697
Magic Gamma
Telescope
0.1875 1552.0641 1552.2750 1552.7039
Rice Cammeo
And Osmancik
211.2688 358.2641 378.0189 297.6197
Wine 21.0234 40.3781 43.8195 73.9416
Abalone 0.1875 294.4004 294.5879 294.9893
Air Quality 880.1270 1294.5430 1373.1387 1233.9355
Appliances
Energy
Prediction
4164.2959 5894.4902 6261.5137 5832.6855
Concrete
Compressive
Strength
65.6523 137.7617 144.9668 94.3340
Forest Fires 43.2832 87.1992 92.4219 157.8330
Real Estate
Valuation
22.2676 43.1992 46.3398 79.0088
Wine Quality 0.1875 624.3691 624.5879 625.0762
V Limitation

### v4

@@ V Results & Discussion
This section presents the empirical results of our extensive experiments. We selected 5 representative models and 3 scaling techniques, along with a baseline without scaling, as this subset already allows us to draw meaningful conclusions and observe distinctions among models. The complete results are available in the Appendix A.
V-A Impact on Validation Metrics
We used the Wilcoxon signed-rank test[119, 120] to compare the performance of the algorithms on scaled and unscaled datasets and the Friedman test[121] to determine the statistical significance of the data scaling results. We adopted a significance threshold of 0.01 to reduce the likelihood of Type I errors (false positives) when comparing models with and without feature scaling. Given the multiple datasets and models involved, a stricter alpha level provides greater confidence that any observed differences in performance are truly meaningful. The complete results of these statistical tests are in Appendix A.
As anticipated, one of the key findings of our classification experiments is the differential impact of feature scaling on model performance. Ensemble methods, including Random Forest and the gradient boosting family (LightGBM, CatBoost, XGBoost), demonstrated strong robustness by consistently achieving high validation performance regardless of the preprocessing strategy or dataset. The Naive Bayes model showed similar scaling resistance, although its overall accuracy was not competitive with these top-tier ensembles. This observation is statistically supported by the Wilcoxon signed-rank test comparing scaled versus unscaled features. These results are confirmed by an analysis of Figure 2 and Tables IV, LABEL:tab:accuracy_stests, and LABEL:tab:accuracy_big_table.
In contrast, the performance of LR, SVM, KNN, TabNet, and MLP were highly dependent on the choice of scaler, revealing their pronounced sensitivity to data preprocessing (see Figure 2 and Table LABEL:tab:accuracy_big_table). This observation is statistically supported by the Wilcoxon signed-rank test comparing scaled versus unscaled features, as shown in Tables IV, and LABEL:tab:accuracy_stests.
Similarly, the Friedman test across all scaling methods confirmed significant performance variation for LR, SVM, MLP, and KNN, while RF, NB, CART, AdaBoost, CatBoost, LGBM, XGBoost, and TabNet showed no significant differences. These results quantitatively support the observed trends, highlighting which models are inherently scaling-sensitive versus scaling-robust.
Figure 2: Accuracy results for the 9 datasets. The ensemble models are: RF, LGBM, CatBoost, and XGBoost.
In regression tasks, a similar pattern of scaling sensitivity was observed when employing the regression counterparts of these classification models (see Figures 3, 4, and 5), with the results from the Wilcoxon signed-rank and Friedman tests shown in Tables IV, LABEL:tab:r2_stests, LABEL:tab:mse_stests, and LABEL:tab:mae_stests. Notably, for KNN in regression, the Wilcoxon test did not indicate statistically significant differences across scaling methods. This suggests that the underlying mathematical principles governing these model families lead to consistent behavior regarding data scaling, whether applied to classification or regression problems. The complete results are presented in Tables LABEL:tab:r2_big_table, LABEL:tab:mse_big_table and LABEL:tab:mae_big_table.
Figure 3: MSE results for the 7 datasets. The ensemble models are: RF, LGBM, CatBoost, and XGBoost.
TABLE IV: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for classification tasks and regression tasks.
Model Wilcoxon $p$ Wilcoxon sig. Friedman $p$ Friedman sig.
SVM 0.0001 Yes 0 Yes
SVR 0 Yes 0.0007 Yes
KNN 0-0.6811 Yes/No 0.0393-0.0040 Yes/No
MLP 0-0.0077 Yes 0-0.0006 Yes
Ensemble 0.0252-0.9710 No 0.1369-1.0 No
@@ V-A Impact on Validation Metrics
V-B Impact on Training and Inference Times
The application of different scaling techniques had a variable impact on inference times across the evaluated models. Notably, Classification and Regression Trees based models exhibited exceptionally robust behavior, remaining unaffected across all scaling methods and datasets. While certain Machine Learning algorithms, such as K-Nearest Neighbors (KNN), Support Vector Machine (SVM), and Support Vector Regressor (SVR) showed more evident sensitivity to the choice of scaling technique, as illustrated in Figures 7 and 9, this was not the norm. As shown in Tables LABEL:tab:inf_class_big_table and LABEL:tab:inf_reg_big_table, for the majority of models, the preprocessing step introduced only a small, non-uniform computational overhead, which may become more significant in the context of large-scale datasets or time-sensitive applications requiring real-time inference.
For training time, the effects of scaling largely mirrored those seen in the validation accuracy results, as shown in Figures 6 and 8. Certain models, notably tree-based ensembles, did not derive a significant speed benefit from feature scaling, while others were more sensitive. The complete results are available in Tables LABEL:tab:train_class_big_table and LABEL:tab:train_reg_big_table.
V-C Results of Memory Usage (kB)
The analysis of Figure 10 reveals a clear distinction in memory consumption between scaling techniques. As expected, applying no scaling resulted in zero additional memory consumption, beyond what is required to load the data. Among the actual scaling methods, the RobustScaler, StandardScaler, Tanh Transformer, and Hyperbolic Tangent were found to be the most memory intensive. In contrast, the MaxAbsScaler, MinMaxScaler, and Decimal Scaler consistently registered the lowest memory usage, as detailed in Table LABEL:tab:memory_big_table.
VI Limitation